OpenStudy (anonymous):

Which of the following ordered pairs represents the solution to the system given below? 2x + y = 20 4x − 2y = 40 A) (10,0) B) (0,10) C) (4,2) D) (10, -10)

OpenStudy (kidrah69):

We can use elimination to solve the problem :) do you know what the first step would be?

OpenStudy (anonymous):

no @kidrah69

OpenStudy (kidrah69):

Alright :) First we need to get opposite signs! \[-2(2x + y = 20)\] distribute -2 into everything in the parentheses

OpenStudy (kidrah69):

-2 * 2x=? -2*y=? -2*20=?

OpenStudy (anonymous):

-4 ??? -40 @kidrah69

OpenStudy (kidrah69):

1) -4x 2)-2y 3) -40 Moving on... \[ \frac{ -4x-2y=-40 }{ 4x − 2y = 40 } \]

OpenStudy (kidrah69):

-4-4=? -2-2=? -40+40=?

OpenStudy (anonymous):

-8 -4 0

OpenStudy (kidrah69):

My mistake sorry -4+4=0 ok so we have -4x=0 -4 divided by 0=?

OpenStudy (anonymous):

0

OpenStudy (kidrah69):

y value is 0

OpenStudy (kidrah69):

Without doing any extra work what is the answer? :)

OpenStudy (anonymous):

zero ?

OpenStudy (anonymous):

@kidrah69